OpenStudy (anonymous):

(cos(x)^3- sin(x)^3)/cos(x) - sin(x)

OpenStudy (anonymous):

factor numerator as difference of two cubes

OpenStudy (saifoo.khan):

(cos(x)^2- sin(x)^2)

OpenStudy (anonymous):

\[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (saifoo.khan):

am i right satellite?

OpenStudy (anonymous):

ok so how do I solve it after that

OpenStudy (anonymous):

in your case you have \[\frac{(\cos(x)-\sin(x))(\cos^2(x) +\cos(x)\sin(x)+\sin^2(x)}{\cos(x)-\sin(x)}\]

OpenStudy (anonymous):

cancel to get \[\cos^2(x)+\cos(x)\sin(x)+\sin^2(x)\]

OpenStudy (anonymous):

and then use \[\cos^2(x)+\sin^2(x)=1\] to get \[1+\cos(x)\sin(x)\]

OpenStudy (anonymous):

can I ask you anoth question

OpenStudy (anonymous):

go ahead